(9x+2)+(4x-1)-(6x+3)+4(x-1)=12
confused with the previous problem and answer. New one. thanks so much

Found 2 solutions by stanbon, asha:
Answer by stanbon(75887)

(9x+2)+(4x-1)-(6x+3)+4(x-1)=12
9x+4x-6x+4x+2-1-3-4=12
11x-6=12
11x=18
x=18/11
Cheers,
Stan H.

Answer by asha(30)

(9x+2)+(4x-1)-(6x+3)+4(x-1)=12
9x+2+4x-1-6x-3+4x-4=12
take all x terms to the l.h.s and number terms to the r.h.s.
9x+4x-6x+4x =-2 +1 +3 +4 +12
11x = 18
x=18 /11 is the soln.
=1 7/11 is the soln.